Firm D has COGS equal to $ 500,000 for year ended 12/31/11. Gross Profit for the year also was $ 500,000, and the Profit Margin for the year was 15%. The Retention Rate for the year was 80%. and year-end Assets totaled $ 2,000,000. Year-End Debt was 50% of the Year-End Asset level, and Year 2004 was the firm's first year of operations. Based upon the information, please provide the Statement of Owners Equity for year ended 12/31/11 for Firm D.
Answer #1

Computation of Net Income

Net income = (COGS + Gross Profit) * Profit Margin

Net income = (500000 + 500000) * 15%

Net income = 150000

Closing Retained Earnings = Net Income - Dividend

Closing Retained Earnings = 150000 - 150000 * (1 - 0.80)

Closing Retained Earnings = $120000

Shareholders' Equity = Total Assets * (1 - Debt Ratio)

Shareholders' Equity = 2000000 * (1 - 0.50)

Shareholders' Equity = $1000000